OpenStudy (anonymous):

Phase Shift for this equation? y=sin(pi x + pi/2)

OpenStudy (anonymous):

is pi/2 right ??

OpenStudy (anonymous):

See if this helps http://www.wolframalpha.com/input/?i=y%3Dsin%28pi+x+%2B+pi%2F2%29

OpenStudy (anonymous):

y=sin(pi x + pi/2) y=sin(3pi x / 2)

OpenStudy (anonymous):

or y=sin(3/2 * pi x)

OpenStudy (anonymous):

satellite please make it short lol >.<!

OpenStudy (anonymous):

lol jk jk :D

OpenStudy (anonymous):

instead of memorizing formula, you can do this as follows. you know the period of sine is \(2\pi\) and every thing happens over the interval \([0,2\pi]\) so you can set \[\pi x+\frac{\pi}{2}=0\] solve for x and then set \[\pi x+\frac{\pi}{2}=2\pi\] solve for x and get period and phase shift that way on the other hand if you want a formula you can write \[\sin(\pi x +\frac{\pi}{2}=\sin(\pi x+\frac{1}{2})\] and see that the period is now 2 and the phase shift is one half

OpenStudy (anonymous):

did u look at the website @Ala123

OpenStudy (anonymous):

\[\sin(\pi x +\frac{\pi}{2})=\sin(\pi (x+\frac{1}{2}))\]

OpenStudy (anonymous):

yeah i looked but im still confused @@

OpenStudy (anonymous):

good idea. you will see that you have a function that looks just like cosine, but with a period of 2 instead of \(2\pi\)

OpenStudy (anonymous):

yeah the period and amplitude i got it, but the phase shift....btw the answer for phase shift is -1/2 ? why u have 1/2 ?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@satellite73 why the answer is -1/2 ? :|